The sum of the values of x satisfying the equation 233x–2 + 211x+2 = 222x+1 + 1 is

Grade:12

2 Answers

Vikas TU
4 years ago
Dear student 
 233x–2 + 211x+2 -222x+1  =
Take log on both side 
33x-2 + 11x +2 -22x-1 = 0 
22x = 1 
x = 1/22 
Hope this helps 
 
 
Aditya Gupta
4 years ago
once again, vikas’s answer is the most stupid, childish and wrong answer you can expect. he doesnt even know how to use logarithms.
correct method is as follows:
233x–2 + 211x+2 = 222x+1 + 1
multiply with 2^2 both sides
233x + 211x+4 = 222x+3 + 4
or 233x + 211x*24 = 222x*23 + 4
now, let y= 211so that y^2= (211)^2= 211x*2 222x  and y^3= 2^33x
so we have y^3 + 16y= 8y^2 + 4.
you can easily solve this polynomial eqn and you ll get 3 roots y0, y1 and y2. reject the negtive roots (if any).
the value of x would then be [log2(y)]/11.
kindly approve :)